OpenStudy (lalaland_lauren):

Organize the following polynomial expressions from least to greatest based on their degree: x + 2xyz 9x3y2 18x2 + 5ab − 6y 4x4 + 3x2 − x − 4

OpenStudy (anonymous):

\[x + 2xyz=x + 2x^1y^1z^1\] has degree \(1+1+1=3\)

OpenStudy (anonymous):

\[9x^3y^2\] has degree \(3+2=5\)

OpenStudy (anonymous):

i bet you can do the rest

OpenStudy (anonymous):

if not, i can check

OpenStudy (lalaland_lauren):

would this be number three? \[18^2 +5a^1b^1-6y^-1\]

OpenStudy (lalaland_lauren):

that's -6y^1 but I think the one is negative

OpenStudy (anonymous):

\[18x^2 + 5ab − 6y\] has degree 2

OpenStudy (lalaland_lauren):

oh okay so if it doesn't show an exponent then you kind of make your own

OpenStudy (anonymous):

yeah i have a feeling you don't quite get this am i right?

OpenStudy (lalaland_lauren):

yup

OpenStudy (anonymous):

ok here is a polynomial with three terms \[xy^2+x^2y^2+2xy\]

OpenStudy (anonymous):

each term has a degree ' the degree of \[xy^2\] is \[1+2=3\] th degree of \(x^2y^2\) is \(2+2=4\) and the degree of \(2xy\) is \(1+1=2\)

OpenStudy (lalaland_lauren):

so the degrees are the exponents added up

OpenStudy (anonymous):

the "degree of the polynomial" is the degree of the term of highest degree, so the degree of \[xy^2+x^2y^2+2xy\] is \(4\)

OpenStudy (anonymous):

yes, the degree of each term is the sum of the degrees of each variable so for example the degree of \(xy^2z^4\) is \(7\)

OpenStudy (lalaland_lauren):

and whichever has the highest degree is the degree of the polynomial?

OpenStudy (lalaland_lauren):

sorry if I repeat things, just trying to get it through to myself

OpenStudy (anonymous):

that is fine

OpenStudy (anonymous):

you are right, the degree of the polynomial is the degree of the term of highest degree

OpenStudy (anonymous):

so for example if you just have on variable, the degree of say \[2x^5+3x^2+5x+1\] is \(5\) but if you have more than one variable you have to add the degrees of each variable

OpenStudy (anonymous):

with that in mind, what is the degree of \[4x^4 + 3x^2 − x − 4\]?

OpenStudy (lalaland_lauren):

the degree of the polynomial would be 4 because of 4x^4

OpenStudy (anonymous):

yes

OpenStudy (lalaland_lauren):

\[4x^4+3x^2-x-4\] first set has a degree of four, second set has a degree of two, and the third I think it's one but I'm not sure if it's negative one

OpenStudy (anonymous):

a polynomial cannot have a term of negative degree

OpenStudy (lalaland_lauren):

x + 2xyz degree of 3 9x3y2 degree of 5 18x2 + 5ab − 6y degree of 2 I'm pretty sure 4x4 + 3x2 − x − 4 4

OpenStudy (anonymous):

yes

OpenStudy (lalaland_lauren):

okay gotcha

OpenStudy (anonymous):

btw the degree of a constant is zero

OpenStudy (lalaland_lauren):

III, I, IV, II IV, I, II, III III, II, IV, I IV, III, I, II these were the answer choices btw lol

OpenStudy (lalaland_lauren):

what's a constant?

OpenStudy (anonymous):

lol i will let you order them yourself, you got them all right

OpenStudy (lalaland_lauren):

it's a

OpenStudy (anonymous):

\[x^2-5x+\color{red}3\] the constant is \(\color{red}3\)

OpenStudy (anonymous):

yeah it is A

OpenStudy (lalaland_lauren):

so one with no exponent or variable, just a flat out number

OpenStudy (anonymous):

right called a "constant" because it does not depend of what \(x\) is

OpenStudy (anonymous):

i have a question do you know what \\[2x^3+5x^2+4x+7\] is if \(x=10\)?

OpenStudy (lalaland_lauren):

\[2(10)^3+5(10)^2+4(10)+7\] \[2(1,000)+5(100)+40+7\] two thousand five hundred and forty seven

OpenStudy (anonymous):

right so you see we write our whole numbers as polynomials, using base ten (instead of x)

OpenStudy (anonymous):

\[2(10)^3+5(10)^2+4(10)+7=2547\]

OpenStudy (anonymous):

keep that in mind when you learn to add, subtract and multiply polynomials it works pretty much the same way as with whole numbers of course there are some differences, but it is basically the same

OpenStudy (lalaland_lauren):

okay :-) can you help me with some more questions?

OpenStudy (lalaland_lauren):

it's like two or three more, one I think I got down

OpenStudy (anonymous):

sure

OpenStudy (lalaland_lauren):

4x^2 + 3xy + 12yz this would be a second degree trinomial correct?

OpenStudy (anonymous):

yes

OpenStudy (lalaland_lauren):

okay that's what I put, next one

OpenStudy (lalaland_lauren):

Organize the following expressions from greatest to least by number of terms: x + 2xyz 9x2yz 18x2 + 5ab − 6y 4x3 + 3x2 − x − 4 Answers: III, IV, I, II IV, I, II, III IV, III, I, II III, II, IV, I

OpenStudy (lalaland_lauren):

Is it C?

OpenStudy (lalaland_lauren):

Because I think that terms are the sets in the polynomial

OpenStudy (anonymous):

the "number of terms"? yeah we can count for sure

OpenStudy (anonymous):

you are right, it is C

OpenStudy (lalaland_lauren):

Which statement best demonstrates why the following is a non-example of a polynomial? \[\frac{ 33y^2 }{ x^2 } - 62y^2xz-35z^2y^2\] The expression has a variable raised to a negative exponent. The expression has a negative coefficient. The expression has a variable raised to a fraction. The expression has a variable in the denominator of a fraction.

OpenStudy (lalaland_lauren):

I feel like it's C again

OpenStudy (anonymous):

not this time

OpenStudy (anonymous):

a polynomial cannot have a variable in the denominator

OpenStudy (anonymous):

" The expression has a variable raised to a fraction." actually makes no sense at all go with D

OpenStudy (lalaland_lauren):

oh okay, I thought fractions weren't supposed to be in polynomials period lol

OpenStudy (anonymous):

unless it means something iike \[x^{\frac{2}{3}}\]which is NOT a polynomial

OpenStudy (anonymous):

the coeficients can be fractions \[\frac{4}{5}x^2\] is a polynomial

OpenStudy (lalaland_lauren):

Last one :) Which of the following shows 9x^2y − 4x + 3y^3x − 2y^2 written in standard form? 9x^2y − 4x + 3y^3x − 2y^2 3y^3x − 2y^2 + 9x^2y − 4x 9x^2y − 4x − 2y^2 + 3y^3x 3y^3x + 9x^2y − 2y^2 − 4x

OpenStudy (anonymous):

standard form means order the terms from highest degree to lowest degree

OpenStudy (lalaland_lauren):

Then D

OpenStudy (anonymous):

yes

OpenStudy (lalaland_lauren):

I got a one hundred, thank you so much!!

OpenStudy (anonymous):

yay yw see you in lala land

OpenStudy (lalaland_lauren):

Lol! I'll probably need help tomorrow, will you still be on?

OpenStudy (anonymous):

always glad to help someone who is actually trying

OpenStudy (anonymous):

probably you can always tag me if you like

OpenStudy (lalaland_lauren):

Okay :) thank you so much, I have to go to sleep now. Good night!

OpenStudy (anonymous):

sometime around 9:30 or 10 usually gnight